OpenStudy (anonymous):

simplify (x^-3x-3)+(2x^+7x-2)

OpenStudy (anonymous):

Dont let the parenthesis confuse you, they are meaningless in this statement, (because there is no negative to distribute)

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

i got x^+2x^

OpenStudy (anonymous):

@Werds96 can you try to post it using equation form?

OpenStudy (anonymous):

3x^2+4x-5, now figure out what i did :)

OpenStudy (anonymous):

well i just need to know how to simplify

OpenStudy (anonymous):

your combining like terms here, the parenthesis are there to show order of operations, but because there is a + sign you can get rid of them. (if there was a negative sign then you would have to distribute the negative, switching all the signs within the parenthesis on the right, dont ever forget to distribute a negative!)

OpenStudy (anonymous):

what does mean ^?

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

\[(x^2-3x-3)+(2x^2+7x-2) =x^2+2x^2+7x-3x-2-3=3x^2+4x-5\]